OpenStudy (anonymous):

can someone help me figure out square root of 36 over 5

OpenStudy (anonymous):

Basically, separately find the square root of 36 and the square root of 5

OpenStudy (anonymous):

so the answer would be?

OpenStudy (anonymous):

\[\frac{ \sqrt{36} }{ 5 } or \sqrt{\frac{ 36 }{ 5 }}\] ?? which one?

OpenStudy (anonymous):

36 sqrt over 5sqrt

OpenStudy (anonymous):

separate both of them \[\frac{ \sqrt{36} }{\sqrt 5 }\] what is \[\sqrt{36}\]?

OpenStudy (anonymous):

the square root of 36 is 6

OpenStudy (anonymous):

so \[\frac{ 6 }{ \sqrt{5} }\]...but don't leave the radical at the bottom get rid of it by \[\frac{ 6 }{ \sqrt{5} } x \frac{ \sqrt{5} }{ \sqrt{5} }\]

OpenStudy (anonymous):

your final answer should be \[\frac{ 6\sqrt{5} }{ 5 }\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

YW